(200x)/((10^2+x^2)^1/2)-160=0
need a step by step solution i know the answer i just can't solve it the answer is 40/3

Answer by richwmiller(17219)

(200x)/((10^2+x^2)^1/2)-160=0
(200x)/((100+x^2)^1/2)=160
5x/((100+x^2)^1/2)=4
5x=4*((100+x^2)^1/2)
5/4x=((100+x^2)^1/2)
25/16x=100+x^2
9/16x^2=100
3/4x=10
x=4/3*10
x=40/3